### v2

# Tournament design: A review from an operational research perspective

Abstract
Every sport needs rules. Tournament design refers to the rules that determine how a tournament, a series of games between a number of competitors, is organized. This study aims to provide an overview of the tournament design literature from the perspective of operational research. Three important design criteria are discussed: efficacy, fairness, and attractiveness. Our survey classifies the papers discussing these properties according to the main components of tournament design: format, seeding, draw, scheduling, and ranking. We also outline several open questions and promising directions for future research.
Keywords: OR in sports; ranking; scheduling; seeding; tournament design
1 Introduction
Practicing sports is an important part of the life of many people all over the world as many health, psychological, and social benefits are associated with physical activity. Watching and following sports is even more popular than playing it. About 3 billion people watched the 2020 Summer Olympic Games (Statista, , 2023), while no less than 1.5 billion people tuned in for the 2022 FIFA World Cup final (FIFA, 2023b, ). Even tournament design aspects can be popular: the draw for the 2021/22 UEFA Champions League Round of 16 has attracted more than 1 million views on the YouTube channel of UEFA (Boczoń and Wilson, , 2023). Hence, “designing an optimal contest is both a matter of significant financial concern for the organizers, participating individuals, and teams, and a matter of consuming personal interest for millions of fans(Szymanski, , 2003, p. 1137).
Every tournament is governed by a set of rules that determine the set of contestants, the format, the schedule, the ranking, and the prize allocation. Deciding on these rules is what we call tournament design. In the following, the contestants will be called players or teams interchangeably. The set of matches to be played is given by the tournament format. Seeding is typically used to guarantee that a strong contestant is not eliminated at the early stages of the tournament merely because it plays against an even stronger opponent. Subject to these and other constraints, the contestants can be allocated randomly, through a so-called draw. At first sight, the schedule, the order and timing of the matches, seems a purely operational issue (involving travel distance, referee and venue availability, etc.). However, it may have powerful effects on the fairness and the outcome of the tournament. Ranking aims to order the teams according to their performance in the tournament. Although performance can be directly measured in some sports (e.g. by finishing times), most sports allow only to rank the contestants by paired comparisons, games or matches. Prize allocation determines how prizes (e.g. promotion, relegation, qualification, broadcasting revenue) will be awarded to the contestants.
Given the variety of tournament designs, the question arises on what basis they can be compared. We identify four different criteria that appear frequently in the literature on tournament design: 1) Efficacy; 2) Fairness; 3) Attractiveness; and 4) Strategy-proofness. Efficacy is defined as the ability of a tournament to reveal the “true” (but hidden) ranking of participants. In contrast, fairness has no uniform definition (see Section 2.3.2 and Baumer, (2024)); in our paper, it mainly refers to the principle of equal treatment. Attractiveness also has multiple interpretations, but they are always connected to the excitement that the tournament generates. Finally, all participants should have an incentive to exert full effort and play according to their true strength under any possible scenario in the tournament. That is, the tournament design needs to be strategy-proof.
Our focus on these criteria may suggest that sports federations can freely choose an optimal tournament design from their own perspective, i.e., a design showing an appropriate balance of theoretical properties. In practice, however, traditions and the nature of the sport may play an even larger role in this decision process. For instance, European football leagues are usually played in a round-robin format (Section 2.1.1), while tennis competitions are organized as a seeded knockout tournament (Section 2.1.2). Similarly, the Swiss system (Section 2.1.3) is mainly used in chess. Sports differ in how much time a player needs to recover between two matches: for instance, baseball teams play almost daily, while World Rugby recommends that at least four clear recovery days are provided after each match. The number of rest days required may depend on the frequency of changes allowed by the rules, too; ice hockey teams can play every second day mainly because they often change lines. Some sports plan their season in order avoid overlap with other popular sports (Pagels, , 2018), while travel costs may force tournaments to be organized in a short time span on one location (Nurmi et al., , 2014). Hence, any “good” tournament design depends on the particular sport, and operational research should take these constraints into account to improve decision making. However, in this paper – apart from a number of specific examples – we will maintain a more generic view.
Tournament design is a research field at the interface between operational research and (sports) economics. Operational research mainly contributes to tournament design by optimizing design choices over certain criteria and aiding decision-making. In sports economics, the main issue is typically how the tournament design may affect the behavior and efforts of the players, through theoretical or empirical research (Medcalfe, , 2024; Palacios-Huerta, , 2023). There is a clear synergy between these two angles: sports economics research can bring new criteria and conditions to be taken into account in optimization (or debunk myths that traditionally played a role in tournament design), and validate the decisions proposed by operational research.
In this paper, we assume an operational research perspective. Therefore, prize allocation will be discussed only marginally, as this is, in our view, related to sports economics rather than to operational research. We refer the reader to Moldovanu and Sela, (2001), Bergantinos and Moreno-Ternero, (2020), and Dietzenbacher and Kondratev, (2023) for excellent discussions on this topic. Furthermore, even though this is certainly one of the most important theoretical properties of tournament design, we have decided not to review strategy-proofness here for the same reason. In addition, including this issue would have substantially increased the length of our survey, and a recent book deals with the problem of incentives from an operational research perspective (Csató, 2021c, ). Analogously, the effect of tournament design on competitive balance, consumer demand, or market structure will be mostly ignored; we refer the interested reader to a seminal economics paper on these aspects (Szymanski, , 2003). Nevertheless, we always try to emphasize practical applications by connecting the topics to real-world tournaments. As such, this survey will hopefully be useful for sports economists, too, and advance collaboration between the two domains. Sports administrators and tournament organizers can also benefit from such an up-to-date review of the tournament design literature.
Previous surveys on operational research and tournament design have so far mainly focused on scheduling, and, in particular, its operational aspects, see Rasmussen and Trick, (2008), Kendall et al., (2010), Ribeiro, (2012), Van Bulck et al., (2020), and Ribeiro et al., 2023b . Self-contained textbooks about sports scheduling are Briskorn, (2008) and Ribeiro et al., 2023a . There are also good – albeit somewhat outdated – overviews of the operational research analysis of sports rules (Wright, , 2009, 2014). Kendall and Lenten, (2017) consider a number of examples from sports where rule changes have had unintended consequences. The recent survey of Lenten and Kendall, (2022) collects academic works that proposed rule changes in sports. We add to the existing surveys by considering a broader set of papers and covering a wider set of design issues and criteria, motivated by recent contributions of operational research in these areas. On the other hand, we offer less by excluding many studies that do not explicitly deal with tournament design but focus on, for example, individual games.
The rest of the paper is organized as follows. Section 2 provides the necessary background and terminology on tournament formats, ranking, and the three design criteria presented above. Investigations of tournament formats, seeding policies, draw rules, scheduling, and ranking methods with respect to these criteria are discussed in Sections 37. Finally, Section LABEL:Sec8 concludes and lists several open questions identified by us.
2 Basic terminology and background
Tournament design uses several, not always well-defined, notions. This section attempts to summarize and clarify them in order to prepare and structure the following discussion. In particular, Section 2.1 presents tournament formats, Section 2.2 overviews ranking, and Section 2.3 defines three criteria that play a crucial role in the evaluation of tournament designs.
2.1 Tournament formats
The format of a tournament indicates the set of matches to be played, without identifying the particular teams. Three fundamental types of tournament formats exist: knockout, round robin, and Swiss system tournaments. They are presented in Sections 2.1.1, 2.1.2, and 2.1.3, respectively. More complex tournament formats can be formed by combining these formats, as briefly discussed in Section 2.1.4.
2.1.1 Round robin
One of the basic competition formats is the round robin tournament (Harary and Moser, , 1966). In a $k$-round robin ($k$RR) tournament, each participant faces each other $k$ times. Most domestic football leagues are $k$RR tournaments (Lasek and Gagolewski, , 2018); for example, the competitions in the top five leagues (England, France, Germany, Italy, Spain) are organized as a 2RR. Since 1RR is typically called “single round robin” and 2RR is called “double round robin”, we will use the notations SRR and DRR, respectively, for these formats.
If not all rounds of a round robin tournament are played, by design (especially if there are many contestants), or because the tournament is aborted (see Section LABEL:Sec721), we call it an incomplete round robin tournament. While this format dates back to at least Cochran, (1971), who discussed it in the context of bridge, it got more attention through the qualifying phase of the 2019/20 CONCACAF Nations League, which was contested by 34 teams in a single group, playing just 4 matches each (two home and two away). UEFA adopted an incomplete round robin structure for its club competitions from the 2024/25 season: the group stage is replaced by an incomplete round robin phase with 36 clubs that play against six or eight different opponents (UEFA, , 2023). Incomplete round robin tournaments can also emerge from unevenly combining round robin tournaments between a subset of the opponents (e.g. teams in the same division) and bipartite round robin tournaments, where each team plays against each team from another group (e.g. teams in different conferences, or interleague play) as is common in North American sports.
2.1.2 Knockout
Another traditional competition format is the knockout tournament. In contrast to round robin tournaments, here a player generally faces only a subset of the other players. In a single knockout (SKO) tournament, players are immediately eliminated if they lose.
In a double knockout (DKO) tournament, one loss is allowed in the tournament. This format is organized as two separate SKOs: the first SKO is called the main tournament, and the second SKO is called the consolation tournament. A player losing in the main tournament is moved to the consolation tournament. However, another loss in the consolation tournament eliminates the player. In essence, a player needs to lose twice in order to be knocked out. If the winner of the main round loses the final match against the winner of the consolation tournament, a rematch is played between them. The final phase of the Russian football cup is a double knockout tournament since the 2022/23 season. A variant of this structure is used, for example, in judo, where the loser of the final in the main tournament fights against the winner of the consolation tournament for the second place. Thus, the winner will always have no losses, the silver medalist will have one loss, and the bronze medalist will have two losses.
In some knockout tournaments, the two players facing each other play more than one match. For instance, in the NBA playoffs, the opponents play as a “best-of-7”: the team that first reaches four wins advances to the next round. If multiple knockout tournaments are played by the same players, the SKO is said to be repeated (David, , 1959) or serial (Lambers et al., , 2024). The results of all subtournaments are aggregated to determine the final winner.
[…]
72635481
[…]
532164
[…]
654321
Figure 1: Different types of knockout tournaments
[…]
2.1.3 Swiss system
The Swiss system is a non-eliminating tournament format where a predetermined number of rounds is played. Contrary to the round robin format, it contains fewer rounds than the $n-1$ required to meet all players in a tournament with $n$ players. Its key characteristic is that opponents are determined dynamically. In the first round, the matches depend on an exogenous ranking of the players. In all subsequent rounds, opponents are matched based on several criteria, but the primary goal is typically to play against an opponent having approximately the same number of points (Sauer et al., , 2024). The Swiss system was first used at a chess tournament in Zürich in 1895, but its variants have recently been applied to schedule other sports competitions, including e-sports tournaments (Dong et al., , 2023).
2.1.4 Multi-stage and hybrid tournaments
Tournaments can be played in multiple stages. Typically, only a subset of the players is allowed to proceed to the next stage, based on their performance in the previous stage. If the tournament formats differ in at least two stages, the tournament is commonly called hybrid. The FIFA World Cup, with a SRR group stage followed by a knockout stage for the best teams, is a well-known example.
2.2 Ranking
The ultimate aim of organizing a tournament is to rank the contestants. Section 2.2.1 introduces ranking in an individual tournament, Section 2.2.2 deals with the aggregation of rankings, and Section 2.2.3 gives a concise overview of global rankings, which rank the players on the basis of several tournaments.
2.2.1 Tournament ranking and tie-breaking
A tournament ranking gives an order of the teams based on their achievements in the tournament. In some sports, performance can be directly measured by distances or finishing times. In the case of paired comparisons, the score for each possible match outcome (win or loss, in some cases tie) needs to be decided. In football, for instance, a win is typically worth 3 points, a draw 1 point, and a loss 0 points, which is the (3,1,0) scoring/ranking system.
While various sophisticated ranking methods exist, the most widely used approach is to sum these scores. Naturally, this method does not take the quality of a win into account, i.e. matches played against a strong and a weak opponent yield the same number of points. This can be a problem especially in incomplete round robin and Swiss system tournaments, as will be discussed in Sections LABEL:Sec721 and LABEL:Sec723, respectively.
Summing the scores resembles counting the number of opponents a player has beaten, which is the only ranking rule satisfying three appealing axioms (Rubinstein, , 1980): anonymity (the ranking method does not depend on the identity of the players), positive responsiveness (a loss does not result in a better rank than a win), and autonomous relative ranking (the relative ranking of two players is not influenced by a match where neither of them is involved in). Henriet, (1985) extends this characterization by allowing for ties, which are worth half of a win. While this method is used in some team sports where ties are allowed such as handball, the (3,1,0) scoring system has no theoretical background according to our knowledge.
Another well-known property from social choice theory is inversion (Chebotarev and Shamis, , 1998): the ranking is reversed if all results are reversed. Consider a SRR with four players $i$, $j$, $k$, $\ell$ where $i$ won against $j$ and $\ell$, $j$ won against $\ell$, while the three matches of $k$ resulted in a draw. This example shows that the (3,1,0) scoring system does not satisfy inversion. Indeed, using the notation ($=$, $>$) for the pairwise majority relation based on the aggregated score, the ranking according to the (3,1,0) method is $i>j>k>\ell$, and the ranking after all results are reversed is $\ell>j>k>i$. Naturally, the (2,1,0) scoring system satisfies inversion; in our example, the original ranking is $i>j=k>\ell$, and the ranking after all results are reversed is $\ell>k=j>i$.
Often a complete ranking of the players is desired. Therefore, ranking methods contain various tie-breaking rules to rank players with the same score. In round robin tournaments, the most popular criteria are the number of matches won, head-to-head results, and superior goal difference (Csató, 2021c, , Chapter 1.3). The effects of choosing a specific hierarchy of tie-breaking rules are often non-negligible; for example, they influence the probability of heteronomous relative ranking, i.e. the relative ranking of two teams depends on the outcome of a match in which neither team was involved (Berker, , 2014).
Knockout tournaments also require tie-breaking rules in some sports such as football and handball. Tied matches are usually followed by extra time, and the final tie-breaking is usually a penalty shootout since the 1970s (Palacios-Huerta, , 2014). UEFA introduced the away goals rule in the 1965/66 European Cup Winners’ Cup for two-legged clashes. According to this principle, if both teams score the same number of goals in the two games, the team that scores more goals away will advance.
The majority of tie-breaking rules used in Swiss-system tournaments are based on the idea that players playing against stronger opponents should be preferred. One of the most extensively used methods is Buchholz, which is essentially the arithmetic mean of the scores obtained by the opponents. Freixas, (2022) suggests using a weighted average of the opponents’ scores where the weights are assigned according to the binomial distribution.
2.2.2 Aggregation of rankings
Some tournaments consist of individual contests, which have their own rankings. Examples include cycling (e.g. Tour de France), motorsport racing (e.g. Formula One), or winter sports (e.g. Biathlon World Cup). Here, a score is typically associated with each position in each contest, and they are added to obtain the final tournament ranking. A similar aggregation of rankings can be favourable even if the official rule is based on the “objective” finishing times: Ausloos, (2024) argues for ranking teams in cycling races by the cumulative sums of the places of their riders instead of their finishing times. However, the aggregation can lead to different paradoxes; Kaiser, (2019) presents some of them from Formula One. Since rank aggregation is an extensively studied problem in social choice theory, we do not go into details; a recent summary is provided by Kondratev et al., (2023).
2.2.3 Global rankings
Ranking can also refer to the global ranking of players in a particular discipline. In that case, the ranking is often based on a rating, which quantifies their strength relative to each other. Since players in this case typically did not face equally strong opponents, the rating should account for the strength of their opponents. The widely used rating method developed by Arpad Elo, a Hungarian-born physics professor, does exactly that (Elo, , 1978; Aldous, , 2017). After each match, the rating of a player is updated by the formula
$r_{new}=r_{old}+K(S-\mu),$ (1)
where $K$ is a sports-specific parameter reflecting the maximum possible adjustment per game, $S$ is the outcome of the match (typically 1 for a win and 0 for a loss), and $\mu$ represents the expected outcome. The latter is determined by the ratings of players $r_{i}$ and $r_{j}$, as well as the scaling parameter $s$ according to the following formula:
$\mu=\frac{1}{1+10^{(r_{j}-r_{i})/s}}.$ (2)
Naturally, the Elo method can be parameterized in different ways depending on the value of $K$, the definition of match outcome $S$, or the adjustment of $r_{i}$ due to the potential home advantage. Gomes de Pinho Zanco et al., (2024) present a comprehensive stochastic analysis of the Elo algorithm in round robin tournaments, and investigate the effect of parameters $K$ and $\mu$ on its performance to obtain design guidelines.
FIFA, for instance, after having received serious criticism in the academic literature (Lasek et al., , 2016; Cea et al., , 2020; Csató, 2021c, ; Kaminski, , 2022), adopted the Elo method for their world ranking (FIFA, , 2018). The popularity of Elo ratings is shown by the fact that the FIFA world ranking has an alternative, the World Football Elo Ratings (https://www.eloratings.net/), that takes home advantage and goal difference into account as suggested by Szczecinski and Roatis, (2022). Analogously, the performance of football clubs is measured by at least two independent websites, http://clubelo.com/ and http://elofootball.com/. Nonetheless, some global rankings do not take the strength of opponents into account. For example, UEFA club coefficients are based on awarding points for wins and reaching certain stages in previous UEFA club competitions (Dagaev and Rudyak, , 2019).
There are many sports-specific ranking rules and rating systems. Stefani, (2011) lists the official rating systems of 159 sports federations, while Langville and Meyer, (2012) provide an overview of the mathematics of various ranking methods. An overview of ranking systems in football can be found in Van Eetvelde and Ley, (2019).
2.3 Design criteria
Sections 2.3.12.3.3 shortly discuss the three main criteria that have been identified to evaluate a tournament design: efficacy, fairness, and attractiveness.
2.3.1 Efficacy
The ability of a tournament to rank contestants according to their true strength is called efficacy (Lasek and Gagolewski, , 2018; Sziklai et al., , 2022). Efficacy with respect to only the best player is also known as effectivity (Glenn, , 1960) or predictive power (Ryvkin and Ortmann, , 2008; Vu and Shoham, , 2011).
By far the most popular way of evaluating these properties is to use simulation: all matches of the tournament are simulated and the observed ranking is compared with the assumed true ranking of players.By aggregating the results over many simulation runs (i.e. Monte Carlo simulation), meaningful insights can be derived about how efficacious the tournament is. Efficacy is mostly studied from the perspective of the tournament format and the seeding rule.
2.3.2 Fairness
In contrast to efficacy, there is no uniform definition of fairness as there exists no universally agreed system of ethics (Baumer, , 2024; Palacios-Huerta, , 2023). Still, we have identified two main principles of fairness. The first one is based on the idea that, a priori, a player should not be disadvantaged compared to any weaker player. This means that, for every player, the probability of winning the tournament should be higher than that of any weaker player. This concept is also commonly called order preservation (Schwenk, , 2000; Groh et al., , 2008; Vu and Shoham, , 2011; Prince et al., , 2013; Karpov, , 2016; Placek, , 2023).
The second approach of fairness takes on a completely different perspective. It is based on the principle of treating all players equally and tries to eliminate any favoritism towards particular players. For instance, if players are allocated to groups based on a draw, the procedure is said to be fair if all allocations are equally likely. Analogously, fairness in scheduling is mostly related to this principle by balancing some objectives (the number of rest days, the distances to be traveled, the number of consecutive home or away games, etc.) over the teams. Equal treatment of all players is sometimes weakened to equal treatment of equally strong players (Guyon, 2018b, ; Arlegi and Dimitrov, , 2020; Csató, 2022a, ).
Note that the first fairness principle is closely related to efficacy. In fact, the main reason why an organizer would want to adhere to this principle is to guarantee efficacy. Hence, to reduce the overlap between the discussion of our design criteria, fairness will mean the principle of equal treatment in the following. Therefore, papers studying fairness from the perspective of order preservation will be presented in the sections dealing with efficacy.
2.3.3 Attractiveness
The attractiveness of a match can be defined in multiple ways: by its quality, competitive intensity, and importance. The quality of a match played by players $i$ and $j$ with their strengths $r_{i}$ and $r_{j}$, respectively, can be quantified as $r_{i}+r_{j}$, while its competitive intensity can be defined as $-\lvert r_{i}-r_{j}\rvert$ (Csató, 2020a, ; Dagaev and Rudyak, , 2019; Dagaev and Suzdaltsev, , 2018). Both measures seem to be linked with higher attendance rates (Forrest and Simmons, , 2002; García and Rodríguez, , 2002) and TV audiences (Schreyer et al., , 2018).
On the other hand, the importance of a match is usually defined with respect to a certain prize. Jennett, (1984) suggested a simple measure to estimate the uncertainty of winning the championship ex post. Let $k$ be the total number of points that turned out to be necessary to win the championship. The importance of a match is the inverse of the minimum number of matches required to achieve $k$ points. This metric was later used by Borland and Lye, (1992) and Dobson and Goddard, (1992) to investigate attendance rates. A decade later, Schilling, (1994) introduced probably the importance measure that is probably most well-known. The Schilling importance ${}_{i}S(X)_{t,t+k}$ of match $t+k$ at time period $t$ for team $i$ is defined as:
${}_{i}S(X)_{t,t+k}=Pr(_{i}X|_{i}W_{t+k},H_{t})-Pr(_{i}X|_{i}L_{t+k},H_{t}),$ (3)
where $Pr(_{i}X|_{i}W_{t+k}(L_{t+k}),H_{t})$ is the probability that team $i$ wins prize $X$ (e.g. avoiding relegation) if $i$ wins (loses) the match $t+k$, given the history $H_{t}$ of all matches played before $t$. Lahvička, 2015b extends Schilling’s measure to account for ties.
Many other ways have been suggested to quantify match importance. Goossens et al., (2012) take into account the number of teams that are still competing for a given prize. An integer program is used to determine the highest and lowest rank any team can still obtain at the end of the season when a certain match is played. Similar formulations are proposed by Ribeiro and Urrutia, (2005); Raack et al., (2014); Gotzes and Hoppmann, (2022). Geenens, (2014) measures the decisiveness of a match based on the expected entropy (i.e. uncertainty). Corona et al., (2017) use an entropy-based measure to identify the most decisive matches in the 2014 FIFA World Cup and the 2015 Copa América. Goller and Heiniger, (2023) extend the entropy measure by considering the difference between the reward probability distributions derived from the possible outcomes of a single event. Csató et al., 2024b introduce a classification scheme to determine weakly (where the final rank of one team is known, regardless of the match outcome) and strongly (where the match outcome does not influence the ranking of both teams) stakeless matches.
There is a wide consensus in the literature that still being in contention for a prize positively affects attendance (García and Rodríguez, , 2009; Pawlowski and Nalbantis, , 2015; Losak and Halpin, , 2024). In particular, the tightness of the contention seems to control the size of this effect (Krautmann et al., , 2011). Furthermore, Buraimo et al., (2022) show that match importance is the key driver for the size of TV audiences in the case of English football games instead of the much more researched outcome uncertainty (see the so-called outcome uncertainty hypothesis (Rottenberg, , 1956), and Schreyer and Ansari, (2022) and van der Burg, (2023) for recent overviews).
While importance is usually measured at the level of individual games, suspense refers to uncertainty in the outcome of the tournament. Although suspense can be related to any outcome, such as relegation or qualification for international competitions, it is typically interpreted within the context of the tournament winner. Suspense is an important measure of attractiveness as it is positively related to both stadium attendance (Pawlowski and Nalbantis, , 2015) and TV audience (Bizzozero et al., , 2016; Buraimo et al., , 2020, 2022).
3 Tournament format
The comparison of tournament formats has received much attention in operational research since determining the format is probably the first task of the contest organizer. Even if the number of games that can be played is strictly limited, usually several options remain available. For example, Guyon, (2020) uncovers a serious drawback of the original format of the 2026 FIFA World Cup and lists 7 alternative options. Rennó-Costa, (2023) presents another one based on double elimination, while Guajardo and Krumer, (2023) develop three further formats. Choosing the best from such a set of feasible designs clearly requires their evaluation with respect to a number of criteria.
3.1 Efficacy
At its inception, the tournament design literature was mainly concerned with the probability of the best player winning under a given tournament format, also called effectivity. Pioneering work was done by David, (1959), who used combinatorial techniques to compute exact winning probabilities for a SKO and SRR. To that end, a series of pairwise win probabilities needs to be assumed; $\pi_{ij}$ is the probability that $i$ beats $j$, which are collected in matrix $\Pi$. David, (1959) assumes a cumulative normal distribution $\Phi(\cdot)$ for the pairwise win probabilities, i.e. $\pi_{ij}=\Phi\{\frac{1}{\sigma}(r_{i}-r_{j})\}$, with $\sigma$ being the variance term and $r_{i}$ being the “true” rating of player $i$. Glenn, (1960) and Searls, (1963) apply a similar approach to evaluate tournament structures, although they assume an arbitrary matrix $\Pi$. All three papers compute the probabilities that other players finish in a certain position as well. Although limited to tournaments with 8 players, the authors find SKO to be the least effective format. Double and serial knockout formats are more effective, in particular, if matches are played as a best of three. For random knockout tournaments, Knuth and Lossers, (1987) compute the chance of player $i$ winning the tournament, assuming that it wins against player $i+1$ with probability $p$ and always defeats player $j$ if $j>i+1$. Adler et al., (2017) derive an upper and lower bound on the probability for of winning the tournament for all players, if player $i$ wins against player $j$ with probability $v_{i}/\left(v_{i}+v_{j}\right)$, where the values $v_{1}$, …, $v_{n}$ are exogenously given.
As the exact computation of these expressions quickly becomes intractable, the standard approach in the literature is to perform Monte Carlo simulation. Using this approach, several papers conclude that round robin tournaments are typically more efficacious than knockout tournaments (Appleton, , 1995; McGarry and Schutz, , 1997; Ryvkin and Ortmann, , 2008; Scarf et al., , 2009; Ryvkin, , 2010; Scarf and Yusof, , 2011; Csató, 2021b, ). Furthermore, the efficacy of a SKO increases by extending it towards a DKO (McGarry and Schutz, , 1997; Stanton and Williams, , 2013). Efficacy can usually be improved by increasing the number of matches (Lasek and Gagolewski, , 2018) but the relationship is not necessarily monotonic for complex hybrid tournaments (Csató, 2021b, ). According to earlier studies (Appleton, , 1995; Dinh et al., , 2020), the Swiss system is rather weak in identifying the true ranking of players. However, Sziklai et al., (2022) find that the Swiss system outperforms other formats with the same number of games. To the best of our knowledge, the efficacy of incomplete round robin tournaments has not been examined yet.
Lahvička, 2015a studies the effect of the additional playoff stage on the efficacy of a DRR in the Czech ice hockey league. The average probability that the strongest team becomes a champion is found to decrease from 48% to 39%. The reduction is even higher if the strongest team is more dominant, and is robust with respect to relaxing the assumptions of constant team strengths and the lack of strategic behavior. This is an important message to administrators: uncertainty can be increased by organizing an extra knockout phase after a round robin tournament.
[…]
Table 1: Studies on the efficacy of tournament formats
(a) Papers focusing on efficacy
Study Format $\pi_{ij}$ Best Worst $n$
[…]
SRR,
[…]
${}_{\text{S}}$, SRR+DKO,
${}_{\text{S}}$, DRR+DKO,
[…]
Scarf and Yusof, (2011) 23 formats Poisson()($\cdot$)( ⋅ ) SRR SKO 32
Stanton and Williams, (2013) SKO, DKO CRM, $\frac{1}{2}+\frac{j-i}{2(N-1)}$ DKO SKO 8–$10^{24}$
[…]
SRR, Swiss, Reaper T NS SRR SKO 8
[…]
D(8+6), D(4×$\times 7$),
D(4×$\times 6$)1 $\frac{(57-r_{i})^{k}}{(57-r_{i})^{k}+(57-r_{j})^{k}}$ D(8+6) D($4\times 6$) 24, 28
[…]
SRR, SKO, G64,
G66, G642 $\frac{1}{1+\left[\left(r_{i}+24\right)/\left(r_{j}+24\right)\right]}$ G66 G64 24
[…]
SKO-3, DKO, SRR+SKO $\frac{1}{1+10^{(r_{j}-r_{i})/400}}$ Swiss SKO 32
(b) Papers focusing on effectivity
Unsurprisingly, there exist unavoidable trade-offs between these fairness criteria. The distance minimization and the break maximization problems are equivalent if the distance between every pair of teams is equal (Urrutia and Ribeiro, , 2006). Craig et al., (2009) focus on the trade-off between minimizing travel distance for teams, avoiding sequences of consecutive home or away games, and ensuring equity in rest times. For a SRR with $n$ teams, the minimal number of breaks is $n-2$, which is achieved (for instance) by the canonical schedule. However, this schedule also yields maximally unbalanced carry-over effects for an even number of teams (Lambrechts et al., , 2018). Günneç and Demir, (2019), Goossens et al., (2020), and Çavdaroğlu and Atan, (2022) discuss the trade-off between minimizing breaks and balancing the carry-over effect. Obviously, there are also trade-offs between fairness and organizational or financial criteria (see e.g. Goossens and Spieksma, (2009); Ribeiro and Urrutia, (2012)).
6.3 Attractiveness
In the sports scheduling literature, basically two approaches exist to deal with attractiveness. One approach assumes that the set of popular matches is known or given by a broadcaster, possibly inspired by studies like Wang et al., (2018); Krumer, (2020); Jakee et al., (2022). The schedule should then assign as many of these games as possible to commercially interesting timeslots, and avoid any overlap between them. This approach has been reported in several practical applications, from e.g. Australian football (Kyngäs et al., , 2017), football (Goossens and Spieksma, , 2009; Ribeiro and Urrutia, , 2012), and volleyball (Cocchi et al., , 2018). Matsumura and Hasuike, (2022) estimate the impact of scheduling decisions on TV audiences for the Japanese professional baseball league by multiple regression analysis to develop a schedule that maximizes the expected number of spectators. Domestic football competitions occasionally shut down for some days due to periods reserved for international matches played by national teams. According to Pérez, (2024), home wins have a lower probability after such an interruption, which leads to greater uncertainty of outcome. Consequently, the set of matches played immediately after “FIFA reserved dates” needs to be chosen strategically.
The other, much less attempted approach tries to find a schedule that is as attractive as possible given a prediction of match results, or robust with respect to a subset of possible outcomes. Yi, (2021) proposes an implementer-adversary approach to dynamically develop round robin schedules that maximize suspense, measured by how late in the season the winner is decided. The algorithm works by generating a set of scenarios (realizations of match outcomes) based on historical data from betting odds. The implementer problem finds a schedule that reveals the winner as late as possible, given the information about match outcomes. Next, the adversary problem finds a new (realistic) scenario under which the generated schedule is the worst possible, and this new scenario is added to the set of scenarios. The algorithm continues until the implementer and adversary solutions converge.
We see at least two avenues for future research on improving the attractiveness of tournament schedules. The first is to include match importance measures (see Section 2.3.3) in order to obtain a schedule that maximizes the total importance of matches over the season. This can be easily done in a SRR with three or four teams; papers focusing on stakeless matches usually recommend an optimal (Chater et al., , 2021; Guyon, , 2020) schedule. Using complete enumeration, Csató et al., 2024b have managed to solve this problem in a DRR with four teams and some scheduling constraints. Given the complexity of these measures, such an endeavor will be challenging computationally for larger tournaments. The second avenue focuses on dynamic scheduling, which means that the schedule is not computed and announced fully at the beginning of the season, but is released gradually during the season. Except for recent applications in the NFL (Bouzarth et al., , 2020), in the FIFA World Cup (Stronka, , 2024), and in e-sports (Dong et al., , 2023), dynamic scheduling has not received much attention in the literature. Yet, it may have a great potential to improve the attractiveness of tournaments, as recent information about the strength and ranking of players can be taken into account in order to maximize the importance of the remaining matches.
7 Ranking methods
The theoretical properties of ranking procedures have been widely discussed in the literature. González-Díaz et al., (2014) provide a comprehensive axiomatic analysis of ranking methods for general tournaments where the players may play an arbitrary number of matches against each other. Applying maximum likelihood to the (Zermelo–)Bradley–Terry model (Zermelo, , 1929; Bradley and Terry, , 1952), and the generalized row sum, a parametric family of ranking methods introduced by Chebotarev, (1994), turn out to be the most appealing by satisfying all reasonable axioms that are compatible with each other.
7.1 Efficacy
While efficacy in the context of tournament formats and seeding rules refers to the ability to rank the players according to their true (but hidden) strength, the efficacy of a ranking method usually refers to its predictive power, the ability of a ranking to accurately predict future matches. McHale and Davies, (2007) test a previous formula of the FIFA World Ranking from this perspective. Lasek et al., (2013) compare the predictive power of several methods for ranking national football teams. The Elo rating is found to be a better indicator of success than the old FIFA World Ranking, a result that is reinforced by Gásquez and Royuela, (2016). Indeed, FIFA reformed the World Ranking in 2018 (FIFA, , 2018). Nonetheless, the predictive power of the current FIFA World Ranking can still be considerably improved by incorporating home-field advantage and introducing a weighting by goal difference (Szczecinski and Roatis, , 2022). Somewhat analogously, a variant of the Elo rating (Football Club Elo Ratings, http://clubelo.com/) robustly outperforms the UEFA club coefficient in terms of predictive accuracy for the UEFA Champions League (Csató, 2024a, ).
Table 3: Papers on the efficacy of ranking methods
Tournament design is a research domain that has a strong link with practice. Practitioners and sports governing bodies will hopefully take an interest in this survey, too. In recent years, the results of academic research have frequently been adopted by sports administrators:
  • FIFA has introduced a better ranking procedure for men’s national teams in 2018 (FIFA, , 2018) (Section 2.2.3);
  • UEFA has modified the knockout bracket of the 2020 and 2024 UEFA European Championships to balance group advantages according to the recommendation of Guyon, 2018b (Section LABEL:Sec422);
  • FIFA has reformed the draw of the 2018 and 2022 FIFA World Cups inspired by Guyon, (2014), see Guyon, 2018a (Section LABEL:Sec52);
  • The proposal of Durán et al., (2017) for a fairer schedule of the South American Qualifiers for the 2018 FIFA World Cup has been unanimously approved by the participants and is currently being used (Section LABEL:Sec62).
Governing bodies in sports can also learn from each other. Both the FIBA Basketball World Cup and the FIFA World Cup aim to maximize geographic diversity in the group stage since maximizing intercontinental matches is useful both for attendance and ranking purposes (Guyon, , 2015). On the other hand, the 2023 World Men’s Handball Championship contained a group with two of the three African teams, another group with two of the four South American teams, while the number of European teams varied between one and three in the eight groups.
The example of the 2022 FIFA World Cup draw warns that it is worth consulting with the wider scientific community before rule changes are implemented. On 22 March 2022, FIFA released the details of this draw procedure, which treated the three placeholders of the playoffs unfairly and threatened the balancedness of the groups as discussed in Section LABEL:Sec51. Even though Csató, 2023b immediately realized the problem, it was clearly impossible to correct the mistake before the draw happened.
Consequently, since governing bodies in sports quite often change the design of their tournaments, usually driven by commercial motives, and some innovative ideas repeatedly emerge in practice, operational research in sports will certainly remain a growing research field for the coming decades.
Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838, by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences, and by the Research Foundation - Flanders under Grant S005521N. We are grateful to András Gyimesi, Sergey Ilyin, and three anonymous reviewers for useful remarks.